# Distributional Finite Element curl div Complexes and Application to Quad Curl Problems

### v3

@@ 5.1. Hybridization
The hybridization of the mixed finite element method (39)-(40) is to find $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h},\boldsymbol{\lambda}_{h})\in{\Sigma}_{k-1,h}^{-1}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{\ell+1,h}\times\Lambda_{k-1,h}$ such that
@@
The hybridized mixed finite element method (61a)-(61b) is well-posed, and the solution $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h})\in{\Sigma}^{\rm tn}_{h}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{\ell+1,h}$ coincides with the mixed finite element method (39)-(40).
@@
By (61b) with $\boldsymbol{v}_{h}=0$, $\boldsymbol{\sigma}_{h}\in{\Sigma}^{\rm tn}_{h}$, thus $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h})$ satisfies (39)-(40). ∎
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$ be the solution of the mixed formulation (33)-(34) and the mixed finite element method (61) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
(64) $\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u}_{h}-\boldsymbol{\lambda}_{h}\right\|_{-1/2,h}\lesssim h^{k}(h|\boldsymbol{\sigma}|_{k}+h|\boldsymbol{u}|_{k}+|{\rm curl\,}\boldsymbol{u}|_{k}),$
@@
5.2. Equivalence to other methods
@@
For the hybridized mixed finite element method (61a)-(61b) of the lowest order $k=1$ and $\ell=0,1$, it is also equivalent to a nonconforming finite element method.
@@
(65) $\int_{e}\boldsymbol{v}\cdot\boldsymbol{t}\,q\,{\rm d}s,$ $\;\quad q\in\mathbb{P}_{\ell}(e)\text{ on each }e\in\mathcal{E}(T),$
@@
$W_{h}:=$$\{\boldsymbol{v}_{h}\in{L}^{2}(\Omega;\mathbb{R}^{3}):\,\boldsymbol{v}_{h}|_{T}\in{\rm grad\,}\mathbb{P}_{\ell+1}(T)\oplus(\boldsymbol{x}-\boldsymbol{x}_{T})\times\mathbb{P}_{1}\left(T;\mathbb{R}^{3}\right)\text{for each }$
@@
(67) $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h})-(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h},{\rm grad\,}\phi_{h})$ $=\langle\boldsymbol{f},I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\rangle,$
(68) $(I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},{\rm grad\,}\psi_{h})$ $=0,$
for all $\boldsymbol{v}_{h}\in W_{h}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. Nonconforming finite element method (67)-(68) is a modification of those in [29, 47] by introducing interpolation operator $I_{h}^{{\rm curl\,}}$.
@@
Nonconforming finite element method (67)-(68) is well-posed.
We will show the equivalence between the hybridized mixed finite element method (61a)-(61b) with $k=1$ and nonconforming finite element method (67)-(68).
@@
Let $(\boldsymbol{w}_{h},\phi_{h})\in W_{h}\times\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ be the solution of the nonconforming finite element method (67)-(68). Then $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},\phi_{h},Q_{\mathcal{F}_{h}}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{w}_{h}))\in{\Sigma}_{0,h}^{-1}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{h}\times\Lambda_{h}$ is the solution of the hybridized mixed finite element method (61a)-(61b) with $k=1$.
@@
Choose $\boldsymbol{v}_{h}\in W_{h}$ such that DoF (65) vanishes, then $I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}=0$. Applying the integration by parts on the left hand side of (67), we get
@@
By the arbitrariness of the DoF (66) for $\boldsymbol{v}_{h}$, we obtain $\llbracket\boldsymbol{n}\times({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h})\boldsymbol{n}\rrbracket_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$, that is ${\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h}\in\Sigma_{h}^{\rm tn}$. For all $\boldsymbol{v}_{h}\in W_{h}$ and $\boldsymbol{\mu}_{h}\in\Lambda_{h}$, we get from (23), the integration by parts, and (67) that
@@
On the side hand, for all $\boldsymbol{\tau}_{h}\in\Sigma_{0,h}^{-1}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$, apply (23), (68) and the integration by parts to get

### v4

@@ 5.1. Hybridization
The hybridization of the mixed finite element method (39a)-(39b) is to find $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h},\boldsymbol{\lambda}_{h})\in{\Sigma}_{k-1,h}^{-1}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{\ell+1,h}\times\Lambda_{k-1,h}$ such that
@@
The hybridized mixed finite element method (60a)-(60b) is well-posed, and the solution $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h})\in{\Sigma}^{\rm tn}_{h}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{\ell+1,h}$ coincides with the mixed finite element method (39a)-(39b).
@@
By (60b) with $\boldsymbol{v}_{h}=0$, $\boldsymbol{\sigma}_{h}\in{\Sigma}^{\rm tn}_{h}$, thus $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\phi_{h})$ satisfies (39a)-(39b). ∎
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (60) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
(63) $\left\|\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u}_{h}-\boldsymbol{\lambda}_{h}\right\|_{-1/2,h}\lesssim h^{k}(h|\boldsymbol{\sigma}|_{k}+h|\boldsymbol{u}|_{k}+|{\rm curl\,}\boldsymbol{u}|_{k}).$
@@
5.2. Connection to other methods
@@
For the hybridized mixed finite element method (60a)-(60b) of the lowest order $k=1$ and $\ell=0,1$, it is also related to a nonconforming finite element method.
@@
(64a) $\int_{e}\boldsymbol{v}\cdot\boldsymbol{t}\,q\,{\rm d}s,$ $\;\quad q\in\mathbb{P}_{\ell}(e)\text{ on each }e\in\mathcal{E}(T),$
@@
$W_{h}:=$$\{\boldsymbol{v}_{h}\in{L}^{2}(\Omega;\mathbb{R}^{3}):\,\boldsymbol{v}_{h}|_{T}\in{\rm grad\,}\mathbb{P}_{\ell+1}(T)\oplus(\boldsymbol{x}-\boldsymbol{x}_{T})\times\mathbb{P}_{1}\left(T;\mathbb{R}^{3}\right),\ \forall~{}T\in\mathcal{T}_{h},$
@@
(65a) $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{v}_{h})-(I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h},{\rm grad\,}\phi_{h})$ $=\langle\boldsymbol{f},I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}\rangle,$
(65b) $(I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},{\rm grad\,}\psi_{h})$ $=0,$
for all $\boldsymbol{v}_{h}\in W_{h}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. Nonconforming finite element method (65a)-(65b) is a modification of those in [30, 48] by introducing interpolation operator $I_{h}^{{\rm curl\,}}$. In other words, we identify the complex that accommodates the non-conforming finite elements constructed in [30, 48] and generalize these elements to arbitrary orders.
@@
Nonconforming finite element method (65a)-(65b) is well-posed.
We will show the equivalence between the hybridized mixed finite element method (60a)-(60b) with $k=1$ and nonconforming finite element method (65a)-(65b).
@@
Let $(\boldsymbol{w}_{h},\phi_{h})\in W_{h}\times\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$ be the solution of the nonconforming finite element method (65a)-(65b). Then $({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h},I_{h}^{{\rm curl\,}}\boldsymbol{w}_{h},\phi_{h},Q_{\mathcal{F}_{h}}(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{w}_{h}))\in{\Sigma}_{0,h}^{-1}\times\mathring{\mathbb{V}}^{{\rm curl\,}}_{h}\times\mathring{\mathbb{V}}^{{\rm grad\,}}_{h}\times\Lambda_{h}$ is the solution of the hybridized mixed finite element method (60a)-(60b) with $k=1$.
@@
Choose $\boldsymbol{v}_{h}\in W_{h}$ such that DoF (64a) vanishes, then $I_{h}^{{\rm curl\,}}\boldsymbol{v}_{h}=0$. Applying the integration by parts on the left hand side of (65a), we get
@@
By the arbitrariness of the DoF (64b) for $\boldsymbol{v}_{h}$, we obtain $\llbracket\boldsymbol{n}\times({\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h})\boldsymbol{n}\rrbracket_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$, that is ${\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}\boldsymbol{w}_{h}\in\Sigma_{h}^{\rm tn}$. For all $\boldsymbol{v}_{h}\in W_{h}$ and $\boldsymbol{\mu}_{h}\in\Lambda_{h}$, we get from (24), the integration by parts, and (65a) that
@@
On the side hand, for all $\boldsymbol{\tau}_{h}\in\Sigma_{0,h}^{-1}$ and $\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$, apply (24), (65b) and the integration by parts to get